# Your AI Bosses Are Still Prejudiced: The Emergence of Stereotypes in LLM-Based Multi-Agent Systems

### v1

@@ Introduction
Workplace discrimination and stereotyping have long been recognized as persistent challenges in organizational settings. Social science research has consistently revealed systematic biases in various aspects of professional life, from hiring practices to workplace interactions. A landmark study by (Bertrand and Mullainathan 2004) demonstrated that resumes with White-sounding names received 50% more interview callbacks than those with African American names, highlighting how implicit biases influence recruitment decisions. Recent research confirms the persistence of such discrimination (Kline, Rose, and Walters 2021), while studies like (Kim, Block, and Nguyen 2019)’s work on workplace microaggressions against Asians underscore how these biases manifest in subtle yet impactful ways across different organizational contexts.
@@
Recent theoretical advances by (Bai, Griffiths, and Fiske 2024) suggest that stereotypes can emerge through mechanisms beyond simple cognitive biases or environmental sampling. Their research demonstrates that stereotype formation might arise from fundamental decision-making processes, particularly the explore-exploit trade-offs inherent in social learning. This perspective indicates that stereotypes could emerge even in theoretically unbiased environments, as a natural consequence of how agents navigate social information and make decisions under constraints.
@@
With the rapid advancement of artificial intelligence technologies over the past decade, many researchers and practitioners have suggested that AI systems might exhibit fewer stereotypes and biases than humans when serving in decision-making and leadership positions (Zuiderwijk, Chen, and Salem 2021; Maalla 2021). This optimistic view stemmed from the assumption that machines, being rational and algorithm-driven agents, would inherently make more objective decisions free from human cognitive biases.
However, as Large Language Models (LLMs) (Radford et al. 2019) have become increasingly prevalent, recent studies have strongly challenged this assumption. (Kotek, Dockum, and Sun 2023) found significant gender stereotypes in LLMs’ occupational associations, while (Bai et al. 2025) discovered that even explicitly unbiased LLMs form problematic associations across multiple social domains. In workplace contexts, (An et al. 2024) demonstrated how these biases manifest in hiring decisions, with LLMs showing systematic discrimination patterns based on candidates’ names. While current research has primarily focused on direct bias in model outputs, less attention has been paid to how these biases might emerge and evolve in multi-agent interactions.
@@
Inspired by (Bai, Griffiths, and Fiske 2024), we designed a controlled multi-agent simulation framework that models workplace interactions through numerically-identified agents operating under identical base conditions. Our methodology employs synchronized task-interaction cycles and compares random versus hierarchical task assignment structures (Figure 1). We then quantified stereotype formation using specialized metrics that capture role associations, group biases, competence attributions, and overall stereotype intensity. Testing across multiple LLM architectures revealed consistent stereotype emergence patterns despite the bias-neutral experimental environment.
@@
Following the warmth-competence framework (Bai, Griffiths, and Fiske 2024), we classified jobs into four categories: warm and competent (e.g., data scientist), cold and competent (e.g., manager), warm and incompetent (e.g., rehabilitation counselor), and cold and incompetent (e.g., janitor, truck driver). This classification provides a structured framework for analyzing occupational stereotypes in multi-agent interactions.
@@
The RSI measures the degree to which agents associate specific roles with predefined social categories, where $C_{max}$ denotes the highest category score, $C_{total}$ represents the total score across all categories, and $N$ is the number of available categories. This index captures the strength of role-based associations while accounting for the diversity of available categories through the logarithmic term (Fiske, Cuddy, and Glick 2007).
@@
The GBC quantifies the consistency and strength of group-based evaluations, where AR represents the Agreement Ratio (proportion of consistent evaluations) and NE denotes the normalized entropy of the rating distribution. Through its incorporation of entropy, the coefficient accounts for the diversity of evaluations while identifying the formation of shared biases (Cuddy, Fiske, and Glick 2008).
@@
The CAI measures the strength of competence-based discrimination (Mullen, Brown, and Smith 1992), where $H_{avg}$ represents the mean ratings for high-competence jobs (>5>5 on a 10-point scale), $L_{avg}$ indicates the mean ratings for low-competence jobs (<5<5), and $R_{max}$ denotes the maximum possible rating difference.
@@
To complement our statistical analyses, we implemented a two-stage LLM-based evaluation framework (Chuang et al. 2024). First, an evaluation agent analyzes experimental logs to generate reports on interaction phases, critical events, stereotype development, and group-level effects. Second, a parser agent extracts structured data from these materials, determining stereotype presence and types while identifying specific group effects (Garai et al. 2022). This approach leverages LLMs’ language understanding capabilities while ensuring cross-experimental consistency. The structured outputs from the parser agent are then integrated with our statistical analyses to provide a more comprehensive understanding of stereotype formation patterns in multi-agent interactions.
@@ Conclusion
We will publish a comprehensive GitHub repository contains our complete codebase and experimental data soon.

### v2

@@ Introduction
Workplace discrimination and stereotyping have long been recognized as persistent challenges in organizational settings. Social science research has consistently revealed systematic biases in various aspects of professional life, from hiring practices to workplace interactions. A landmark study by (Bertrand and Mullainathan 2004) demonstrated that resumes with White-sounding names received 50% more interview callbacks than those with African American names, highlighting how implicit biases influence recruitment decisions. Recent research confirms the persistence of such discrimination (Kline et al. 2021), while studies like (Kim et al. 2019)’s work on workplace microaggressions against Asians underscore how these biases manifest in subtle yet impactful ways across different organizational contexts.
@@
Recent theoretical advances by (Bai et al. 2024) suggest that stereotypes can emerge through mechanisms beyond simple cognitive biases or environmental sampling. Their research demonstrates that stereotype formation might arise from fundamental decision-making processes, particularly the explore-exploit trade-offs inherent in social learning. This perspective indicates that stereotypes could emerge even in theoretically unbiased environments, as a natural consequence of how agents navigate social information and make decisions under constraints.
@@
With the rapid advancement of artificial intelligence technologies over the past decade, many researchers and practitioners have suggested that AI systems might exhibit fewer stereotypes and biases than humans when serving in decision-making and leadership positions (Zuiderwijk et al. 2021; Maalla 2021). This optimistic view stemmed from the assumption that machines, being rational and algorithm-driven agents, would inherently make more objective decisions free from human cognitive biases.
However, as Large Language Models (LLMs) (Radford et al. 2019) have become increasingly prevalent, recent studies have strongly challenged this assumption. (Kotek et al. 2023) found significant gender stereotypes in LLMs’ occupational associations, while (Bai et al. 2025) discovered that even explicitly unbiased LLMs form problematic associations across multiple social domains. In workplace contexts, (An et al. 2024) demonstrated how these biases manifest in hiring decisions, with LLMs showing systematic discrimination patterns based on candidates’ names. While current research has primarily focused on direct bias in model outputs, less attention has been paid to how these biases might emerge and evolve in multi-agent interactions.
@@
Inspired by (Bai et al. 2024), we designed a controlled multi-agent simulation framework that models workplace interactions through numerically-identified agents operating under identical base conditions. Our methodology employs synchronized task-interaction cycles and compares random versus hierarchical task assignment structures (Figure 1). We then quantified stereotype formation using specialized metrics that capture role associations, group biases, competence attributions, and overall stereotype intensity. Testing across multiple LLM architectures revealed consistent stereotype emergence patterns despite the bias-neutral experimental environment.
@@
Following the warmth-competence framework (Bai et al. 2024), we classified jobs into four categories: warm and competent (e.g., data scientist), cold and competent (e.g., manager), warm and incompetent (e.g., rehabilitation counselor), and cold and incompetent (e.g., janitor, truck driver). This classification provides a structured framework for analyzing occupational stereotypes in multi-agent interactions.
@@
The RSI measures the degree to which agents associate specific roles with predefined social categories, where $C_{max}$ denotes the highest category score, $C_{total}$ represents the total score across all categories, and $N$ is the number of available categories. This index captures the strength of role-based associations while accounting for the diversity of available categories through the logarithmic term (Fiske et al. 2007).
@@
The GBC quantifies the consistency and strength of group-based evaluations, where AR represents the Agreement Ratio (proportion of consistent evaluations) and NE denotes the normalized entropy of the rating distribution. Through its incorporation of entropy, the coefficient accounts for the diversity of evaluations while identifying the formation of shared biases (Cuddy et al. 2008).
@@
The CAI measures the strength of competence-based discrimination (Mullen et al. 1992), where $H_{avg}$ represents the mean ratings for high-competence jobs (>5>5 on a 10-point scale), $L_{avg}$ indicates the mean ratings for low-competence jobs (<5<5), and $R_{max}$ denotes the maximum possible rating difference.
@@
To complement our statistical analyses, we implemented a two-stage LLM-based evaluation framework (Chuang and others 2024). First, an evaluation agent analyzes experimental logs to generate reports on interaction phases, critical events, stereotype development, and group-level effects. Second, a parser agent extracts structured data from these materials, determining stereotype presence and types while identifying specific group effects (Garai and others 2022). This approach leverages LLMs’ language understanding capabilities while ensuring cross-experimental consistency. The structured outputs from the parser agent are then integrated with our statistical analyses to provide a more comprehensive understanding of stereotype formation patterns in multi-agent interactions.
@@ Conclusion
The code and data of this work are available at https://github.com/advnljs/stereotype.